OpenStudy (tiffany_rhodes):

find partial derivative of f(x,y,z) = xyln(z^2 + xy) with respect to z.

OpenStudy (freckles):

You are to treat everything but z as a constant when taking the partial w.r.t z

OpenStudy (freckles):

So for example do you know how find the derivative of G(z)=2(3)ln(z^2+2(3)) ?

OpenStudy (freckles):

You would use constant multiple rule for the 2(3) part... Chain rule for the ln(z^2+2(3)) part and power rule for the z^2 part

OpenStudy (freckles):

I'm not saying replace x and y with 2 and 3. I'm just saying treat them as you would 2 and 3 or some other constant.

OpenStudy (tiffany_rhodes):

I think my confusion has to do with just taking derivatives of lnx/log(x)

OpenStudy (tiffany_rhodes):

So the derivative of the inside function with respect to z would just be 2z correct?

OpenStudy (freckles):

\[\frac{ d}{dz}\ln(f(z))=\frac{f'(z)}{f(z)}\]

OpenStudy (freckles):

yeah

OpenStudy (tiffany_rhodes):

Because if you treat x*y as a constant, the derivative of a constant is just 0.

OpenStudy (freckles):

correct

OpenStudy (tiffany_rhodes):

So 2z/(xyln(z^2 + xy) ?

OpenStudy (freckles):

the only trouble i have with your answer is the (xy) factor on bottom

OpenStudy (freckles):

I don't like that is it on bottom

OpenStudy (freckles):

\[\frac{\partial f}{\partial z}=xy \cdot \frac{2z}{z^2+xy}\]

OpenStudy (freckles):

the xy in front was just a constant multiple just bring it down

OpenStudy (freckles):

oh and i didn't notice you left the ln intact

OpenStudy (freckles):

that is another problem

OpenStudy (freckles):

\[\frac{ d}{dz}\ln(f(z))=\frac{f'(z)}{f(z)} \] no ln in the derivative here

OpenStudy (freckles):

just differentiating the inside and putting it over the inside

OpenStudy (tiffany_rhodes):

Oh alright. So basically it's just a matter of using the chain rule and knowing the derivative of lnx = 1/x ?

OpenStudy (freckles):

yes

OpenStudy (freckles):

i would stay longer and give you some problems to try but i must go now

OpenStudy (tiffany_rhodes):

No problem, thanks!